OpenStudy (anonymous):

the area of a equilateral triangle ABC is 50cm^2. Find AB. Please show the workout.

OpenStudy (arlen):

area of equilateral triangle= sqrt3/4 * (side)^2

OpenStudy (cwrw238):



OpenStudy (arlen):

\[50^2= \frac{ \sqrt3 }{ 4 }* (side AB)^2\]

OpenStudy (arlen):

@nafis.irteza can you do this?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

is that some kind of formula you wrote?

OpenStudy (arlen):

yes...that is the area of equilateral triangle

OpenStudy (cwrw238):

area = 0.5 * base * height = 0.5*x *h where x = length of each side height h = sqrt(x^2 - (0.5x)^2) 50 = sqrt(x^2 - 0.25x^2) 50 = sqrt(0.75x^2) solve for x x = AB

OpenStudy (arlen):

@cwrw238 i guess AB is a side....so we don't need to do all these stuff....just area of equilateral triangle

OpenStudy (cwrw238):

sorry the equation should be 50 = 0.5 * x * sqrt(0.75x^2) yes - but i think its instructive to see how the formula can be derived

OpenStudy (arlen):

cool :)

OpenStudy (anonymous):

how did u get 0.75? and is sqrt =square root?

OpenStudy (cwrw238):

sqrt = square root x^2 - 0.25x^2 = 0.75x^2 (like 1 - 0.25 = 0.75)

OpenStudy (anonymous):

@cwrw238 how did u figure this out (h = sqrt(x^2 - (0.5x)^2)

OpenStudy (anonymous):

Area of the equilateral triangle = 50 square cm. Area of the equilateral triangle = \[\frac{ \sqrt{3} }{4}*(side)^{2} = 50\] \[(side)^{2} = 50*\frac{ 4 }{ \sqrt{3} }\] \[side = \sqrt{50*\frac{ 4 }{ \sqrt{3} }}\] \[side = \sqrt{200/\sqrt{3}}\] Side = 10.75 approximately.

OpenStudy (anonymous):

@Fall12-13 this really helped...but can u please work it out using trigonometry (sin,cos,tan)

OpenStudy (anonymous):

@Nafis: I don't exactly remember that, sorry :(

OpenStudy (cwrw238):